The height of a triangle is 8m less than the base. The area is 10cm^2. Find the height and the base.
I have tried the equation 10=b(b-8)1/2. I then subtracted 10 from both sides and turned the left side into b^2 -8b-5. Then I could not figure out how to factor correctly. Am I doing the equation wrong or did I mess up on the problem entirely?

Answer by stanbon(75887)

The height of a triangle is 8m less than the base. The area is 10cm^2. Find the height and the base.
I have tried the equation 10=b(b-8)(1/2).
b(b-8) = 20
b^2 - 8b - 20 = 0
----
Factor:
b^2-10b + 2b-20 = 0
b(b-10) + 2(b-10) = 0
(b-10)(b+2) = 0
Positive solution:
base = 10 cm
altitude = b-8 = 2 cm
==========================
Cheers,
Stan H.
==========================